a photon of light a 3.99 x 10-19 J corresponds to radiation in what region and the molecule that aborbs the photon of light would ungergo what transition
same question as above for 1.66 x 10-20 J photon of light
and for 1.99 x 10-25 J
Answer #1

We have the energy of the radiation given here

We have Planck-Einstein relation as ;

E=h\nu

where

E = energy

ν = frequency

h = Plancks's constant = 6.626*10-34 Js

The above equation can be rearranged to find the frequency

\nu =\frac{E}{h}

From the frequency we can refer to the Elecro-magnetic spectrum and find out what region this frequency belongs to.

1.

E = 3.99 x 10-19 J

\nu =\frac{E}{h}=\frac{3.99 \times 10^^^{-19} J}{6.626 \times 10^^^{-34} Js }=6.02\times 10^{14} Hz

Frequency = 6.02*1014 Hz

Radiation with the frequency of the order 6.02x1014 Hz falls in the Visible Region.

Molecule that absorbs photon in the Visible region would undergo electronic transition. Electronic transition is the excitation of outermost of electrons from one orbital to another.

Hence this radiation corresponds  to Visible region and the molecule that absorbs the photon of light would ungergo electronic transition.

2.

E = 1.66 x 10-20 J

\nu =\frac{E}{h}=\frac{1.66 \times 10^^^{-20} J}{6.626 \times 10^^^{-34} Js }=2.51\times 10^{13} Hz

Radiation with the frequency of the order 2.51x1013 Hz falls in the Infrared Region.

Molecule that absorbs photon in the Infrared(IR) region would undergo Vibrational transition. Vibrational transition is the transition of molecule from one vibrational energy level to another.

Hence this radiation corresponds  to Infrared(IR) region and the molecule that absorbs the photon of light would undergo vibrational transition.

3.

E = 1.99 x 10-25 J

\nu =\frac{E}{h}=\frac{1.99 \times 10^^^{-25} J}{6.626 \times 10^^^{-34} Js }=3.00\times 10^{8} Hz

Radiation with the frequency of the order 3.00x108 Hz falls in the Microwave Region.

Molecule that absorbs photon in the Microwave region would undergo Rotational transition. Rotational transition is the transition of molecule from one Rotational energy level to another.

Hence this radiation corresponds  to Microwave region and the molecule that absorbs the photon of light would undergo Rotational transition.